Consider a large clothing shop in Sydney. Suppose it is known that the number of business suits sold per day is normally distributed with a mean, μ = 20 and standard deviation, σ = 10. Mr Wood is employed to sell business suits. The number of business suits he sells each day for 25 days is recorded and the mean number per day calculated.
What is the probability that Mr Wood's average daily sales will be between 15 and 25 business suits? (4 dp)

Answer by ewatrrr(24785)

z = (15-20)/10/sqrt(25) = -5/2
z = (25-20)/10/sqrt(25) = 5/2
P(-5/2 < z < 5/2) = normalcdf(-2.5, 2.5)